Will you please kindly help with this question on compound interest.800 is invested at r% per annum simple interest for 3 years to 168, use formula SI=PRT all divided by 100 to calculate the value of r.
Found 3 solutions by Theo, MathTherapy, ikleyn:
Answer by Theo(13342)

the compound interest formula is f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

the simple interest form is f = p + p * r * n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

the interest part of this formula is i = p * r * n
this makes f = p + p * r * n the same as f = p + i

looks to me like you are interested in solving for i.
the reason you divide r by 100 is because r% / 100 = r.

your problem states that p = 800 and i = 168 after 3 years.
the formula for i becomes:
168 = 800 * r% * 3 / 100
this is the same as:
168 = 800 * r% / 100 * 3 which is the same as:
168 = 800 * r * 3

solve for r to get:
r = 168 / (800 * 3)
solve for r to get:
r = .07

this makes r% equal to r * 100 = 7%.

si = p * r% * t / 100 is the same equation as i = p * r * n
I used i for si and n for t and r for r% / 100.
when p = 800 and r = .07 (7% / 100 = .07) and t = 3, the formula becomes:
i = 800 * .07 * 3 = 168
this is the same as:
si = p * r% * t / 100 where:
p = 800
r% = 7
t = 3
you get:
si = 800 * 7% * 3 / 100 = 168

they didn't really do you any facotrs by explaining the way that they did.
the division by 100 is because r% / 100 = r

let me know if you have any questions regarding this.
keep in mind that compound interest and simple interest are not the same.
this problem was dealing with simple interest only.
theo

Answer by MathTherapy(10552)

Will you please kindly help with this question on compound interest.800 is invested at r% per annum simple interest for 3 years to 168, use formula SI=PRT all divided by 100 to calculate the value of r.
SI = PTR
168 = 800(3)R ----- Substituting 168 for SI, 800 for P, and 3 for T
Interest rate or
It's as SIMPLE as that!

Answer by ikleyn(52781)

.

Let me bring to your mind that the wording/printing in your post is  INCORRECT.

In the first statement, you use the term  "compound interest",

while the second statement is about simple interest.


The account can not be  "simple interest"  and  "compound interest"  at the same time:   banks do not work this way.


A person who posts his  (or her)  posts in section  "Finance",  related to  Finance


                    ===--->>>     MUST  KNOW  IT . . .     <<<---===